OpenStudy (anonymous):

Simplify (x + 2)(x2– 4x – 9)

OpenStudy (anonymous):

\[(x+2)(x^2-4x-9) =x(x^2-4x-9)+2(x^2-4x-9)\]

OpenStudy (johnweldon1993):

For this you would use the *FOIL* method....you take the first term in the first parenthesis (x) and multiply it by EVERY term in the 2nd parenthesis...then you would do the same thing with the 2nd term in the first parenthesis (2) so first...answer these \[\large x \times x^2 = \space ?\] \[\large x \times -4x = \space ?\] \[\large x \times -9 = \space ?\]

OpenStudy (anonymous):

x3– 2x2– 17x – 18 x3– 6x2– x – 18 x3– 4x2– 3x – 18 x3– 10x2– 9x – 18

OpenStudy (johnweldon1993):

And then answer these... \[\large 2 \times x^2 = \space ?\] \[\large 2 \times -4x = \space ?\] \[\large 2 \times -9 = \space ?\]

OpenStudy (anonymous):

these are the answers...

HanAkoSolo (jamierox4ev3r):

@kalynnmarie do you still need help on this question?